# The DESI DR1 Peculiar Velocity Survey: Growth Rate Measurements from the Galaxy Power Spectrum

### v1

@@ 4.1 The models of power spectrum
which essentially performs a Legendre transformation corresponding to Eq.12. The aforementioned theoretical models for the power spectrum can be readily converted into models for the galaxy two-point correlation function, the velocity correlation function, and the galaxy-velocity cross-correlation function using the hankl package [^4] The hankl PYTHON package: https://hankl.readthedocs.io/en/latest/install.html (Karamanis & Beutler, 2021), see Turner et al. (in prep.) and Appendix C for more discussion. The momentum correlation is equivalent to the velocity correlation in linear scales, as demonstrated by Equation 2 of Paper I, therefore, the model of velocity correlation function can be converted from the model of momentum power spectrum approximately.
@@ 4.2 The New Loop Terms for the Cross Power Spectrum
Table 1: The parameters of the power spectrum models. For the density field and the momentum field, each of them have the following set of parameters.
@@ 4.3 The window function convolution
The window function convolution matrix $\bf W$} is constructed using random catalogs, following the approach detailed in Appendix G.2. Each entry of the matrix $\bf W$} is defined according to
@@
For the auto-power spectrum, $w({\bf r})$ and $\bar{n}({\bf r})$ should be computed individually for both the density and momentum fields, and $A^{2}$ is the computed from Eq.7 or 9 respectively. However, when considering the cross-power spectrum, and in accordance with the arguments articulated in Appendix G.2, the following substitutions must be made to the normalization factor in Eq. 48
@@
As depicted in Fig.4, the images illustrates the window function convolution matrix across different types of power spectrum, including the density power spectrum of the BGS data (left panel), the momentum power spectrum of the DESI-PV data (middle panel), and their cross-power spectrum (right panel). Notably, as revealed in Fig.4, the off-diagonal blocks of the $\bf W$} exhibit non-zero values, signifying that the convolution of any given multipole receives contributions from other multipoles. This interdependence implies that higher-order multipoles must be incorporated into the modeling framework, even if they are not directly involved in the final fitting procedure. In this work, we include all multipoles up to $\ell=4$ to ensure comprehensive and accurate representation. We employ $k^{\prime}\in[0,0.4075]\penalty 10000\ h$ Mpc-1 (with 300 bins) in computing each model power spectrum multipole, which should be slightly larger than the $k\in[0,0.3]\penalty 10000\ h$ Mpc-1 which associated with each measured power spectrum multipole, in order to fully cover the diagonal information of each block of the convolution matrix.
@@ 5.1 Fitting method
In this study, we concatenate the density, momentum and cross power spectrum to effectively constrain the growth rate $f\sigma_{8}$. By concatenating the models of these three spectrum, i.e. Eq.29, Eq.30 and Eq.31, we derive a comprehensive framework comprising 13 parameters in total. These encompass the growth rate $f\sigma_{8}$, the biasing parameters $b_{1}\sigma_{8},b_{2}\sigma_{8},b_{s}\sigma_{8},b_{3nl}\sigma_{8}]$ for both the density and momentum fields, as well as the velocity dispersion parameters $\sigma_{vS},\sigma_{vT}]$ for the density and momentum fields, respectively. However, in practical applications, it is not necessary to treat all 13 parameters as independent variables during model fitting with measurements, because these models don’t sensitive equally to every parameter. Through the following testing using mocks, we discover that a streamlined set of only five key parameters $\penalty 10000\ f\sigma_{8},\penalty 10000\ b_{1}\sigma_{8},\penalty 10000\ b_{2}\sigma_{8},\penalty 10000\ \sigma_{vT},\penalty 10000\ \sigma_{vS}\penalty 10000\ ]$ suffices to maintain robustness and precision in the fitting process. This leads us to adopt a unified linear biasing parameter $b_{1}\sigma_{8}$ for both the density and momentum fields, a shared second-order biasing parameter $b_{2}\sigma_{8}$ for both fields. Furthermore, we employ shared velocity dispersion parameters $\sigma_{vT},\penalty 10000\ \sigma_{vS}]$ across both the density and momentum fields.
@@ 5.2 Fitting $f\sigma_{8}$ of the mocks
The MCMC resulting parameter estimates are summarized in Table 2 and illustrated in the right panel of Fig.5. Notably, the estimated value of $f\sigma_{8}=0.468^{+0.061}_{-0.043}$, which closely aligns with the mock fiducial value of $f\sigma_{8,\rm fid}=0.466$ at effective redshift $z_{\rm eff}=0.2$. This agreement demonstrates that our fitting methodology effectively recovers the true growth rate embedded in the mocks, and that the power spectrum models perform robustly up to the non-linear scale of $k_{\text{max}}$=0.3 $h$ Mpc-1 under current conditions. In this analysis, we exclude the momentum power spectrum quadrupole, $P^{p}_{2}$, as it is heavily dominated by noise in survey data. To further assess the adaptability of our model, we also conducted fits incorporating $P^{p}_{2}$ using the mock average, where $P^{p}_{2}$ exhibits sufficient smoothness to yield a reliable fit. For a more detailed exploration, please refer to Appendix E.
@@
Table 2: The MCMC-fit cosmological parameter estimates derived from the BGS and DESI-PV mocks are presented here. These estimates are obtained by fitting the average of 675 BGS mocks and 675 DESI-PV mocks. The degrees of freedom (d.o.f.) are calculated based on 120 data points and 5 free parameters.
@@ 6.1 The systematic error of the FP data
To mitigate this bias, we have to adopt a conservative approach by restricting the momentum power spectrum analysis to $k^{p}_{\text{max}}=0.1\ h\ \text{Mpc}^{-1}$ when fitting $f\sigma_{8}$ from survey data, while retaining $k_{\text{max}}=0.3\ h\ \text{Mpc}^{-1}$ for the density and cross power spectrum. The information encoded in the momentum power spectrum is predominantly concentrated on the larger cosmological scales, specifically at wavenumbers below $k\leq 0.1\ h\ \text{Mpc}^{-1}$ (Howlett, 2019; Bautista et al., in prep.). Consequently, adopting a conservative cutoff at $k^{p}_{\text{max}}=0.1\ h\ \text{Mpc}^{-1}$ enables a robust analysis of the data without sacrificing significant signal content and does not impact our main conclusion.
@@ 6.2 Result of $f\sigma_{8}$
The refined estimate of the growth rate, derived by synthesizing the fitting outcomes from correlation function (Turner et al., in prep.), maximum likelihood estimation (MLE, Lai et al. in prep.) and power spectrum of this paper, yields a consensus value of $f\sigma_{8}(z_{\text{eff}}=0.07)=0.450^{+0.055}_{-0.055}$, in concordance with theoretical predictions too. For a deeper exploration of data synthesis, see Bautista et al. (in prep.).
@@ 6.3 Result of $\gamma$
Table 3: The MCMC-fit cosmological parameter estimates derived from the BGS and DESI-PV data are presented here. The degrees of freedom (d.o.f.) are calculated based on 98 data points and 5 free parameters.
@@ 8 Conclusions
There are several promising avenues for future research building upon this work: (1) Applying our methodology to the Data Release 2 of DESI peculiar velocity survey, which is significantly more homogeneous and extensive than the Data Release 1, thereby enabling a more precise constraint on $f\sigma_{8}$; (2) While the current analysis assumes the local plane-parallel approximation, future efforts will focus on incorporating wide-angle effects into both our models and estimators; (3) As highlighted in Appendix F of this paper, the momentum power spectrum and cross power spectrum lack Galilean invariance, making them susceptible to potential systematic biases in the data. To address this limitation, we aim to transition toward measuring and modeling the pair-wise momentum power spectrum, as proposed by Chen et al. (2025); (4) We have not yet accounted for the gravitational redshift of galaxies. In forthcoming studies, we intend to extend our framework to measure and model this relativistic effect, following the approach outlined by Beutler & Di Dio (2020). The data used to generate the figures in this paper are available at https://zenodo.org/uploads/17672675.
@@ Appendix A Peculiar velocity estimator
where $\sigma_{\eta}$ denotes the measurement error in the log-distance ratio $\eta$. $300^{2}$ km2 s-2 represents the intrinsic scatter in peculiar velocities caused by non-linear galactic motions. As discussed in Appendix A of Paper II, variations in the assumed value of this intrinsic scatter have minimal impact on our results. The $\langle v^{2}({\bf r})\rangle$ term for the DESI-PV random catalog is computed by Bautista et al. (in prep.).
@@ Appendix B Verifying $P^{\delta p}_{\ell}=-P^{\delta p*}_{\ell}$ for cross-power spectrum
The mathematical framework for converting power spectrum to correlation functions primarily draws upon Turner et al. (in prep.). In Fig.14, we present the model correlation functions converted from the power spectrum models presented in Section 4 with parameters from Table 2.
@@ Appendix D The multipoles of the field function
Appendix E Fitting $f\sigma_{8}$ of mocks from $P^{\delta}_{0,2}+P^{p}_{0,2}+P^{\delta p}_{1}$
@@
Table 5: Corresponding to Fig.16, the MCMC-fit cosmological parameter estimates are presented here.
@@
Appendix F The Galilean transformation of power spectrum
@@
Appendix G Deriving the Estimator of Power Spectrum and the Window Funtion Convolution Matrix
@@
G.1 Power Spectrum Estimator

### v2

@@ 4.1 The models of power spectrum
which essentially performs a Legendre transformation corresponding to Eq.12. The aforementioned theoretical models for the power spectrum can be readily converted into models for the galaxy two-point correlation function, the velocity correlation function, and the galaxy-velocity cross-correlation function using the hankl package [^4] The hankl PYTHON package: https://hankl.readthedocs.io/en/latest/install.html (Karamanis & Beutler, 2021), see Turner et al. (2025) and Appendix C for more discussion. The momentum correlation is equivalent to the velocity correlation in linear scales, as demonstrated by Equation 2 of Paper I, therefore, the model of velocity correlation function can be converted from the model of momentum power spectrum approximately.
The PYTHON code for our power spectrum models can be download from the website links provided in Appendix I or see https://github.com/FeiQin-cosmologist/Galaxy_Power_Spectrum/blob/main/CosmPSPy/Code/PSmodFun.py.
@@ 4.2 The new loop terms for the cross power spectrum
Table 1: The parameters of the power spectrum models.  [^6] For the density field and the momentum field, each of them have the following set of parameters.
@@ 4.3 The window function convolution
The window function convolution matrix $\bf W$} is constructed using random catalogs, following the approach detailed in Appendix H.2. Each entry of the matrix $\bf W$} is defined according to
@@
For the auto-power spectrum, $w({\bf r})$ and $\bar{n}({\bf r})$ should be computed individually for both the density and momentum fields, and $A^{2}$ is the computed from Eq.7 or 9 respectively. However, when considering the cross-power spectrum, and in accordance with the arguments articulated in Appendix H.2, the following substitutions must be made to the normalization factor in Eq. 48
@@
As depicted in Fig.4, the images illustrates the window function convolution matrix across different types of power spectrum, including the density power spectrum of the BGS data ( left panel), the momentum power spectrum of the DESI-PV data (middle panel), and their cross-power spectrum ( right panel). Notably, as revealed in Fig.4, the off-diagonal blocks of the $\bf W$} exhibit non-zero values, signifying that the convolution of any given multipole receives contributions from other multipoles. This interdependence implies that higher-order multipoles must be incorporated into the modeling framework, even if they are not directly involved in the final fitting procedure. In this work, we include all multipoles up to $\ell_{max}=4$ to ensure comprehensive and accurate representation. As shown in Fig.5, the differences arising from varying $\ell_{max}$ values are negligible, indicating that $\ell=4$ is sufficiently large to ensure convergence. We employ $k^{\prime}\in[0,0.4075]\penalty 10000\ h$ Mpc-1 (with 300 bins) in computing each model power spectrum multipole, which should be slightly larger than the $k\in[0,0.3]\penalty 10000\ h$ Mpc-1 which associated with each measured power spectrum multipole, in order to fully cover the diagonal information of each block of the convolution matrix.
The PYTHON code for window function convolution can be download from the website links provided in Appendix I or see https://github.com/FeiQin-cosmologist/Galaxy_Power_Spectrum/blob/main/CosmPSPy/Code/PSmodFun.py.
@@
Figure 5: The window function-convolved model power spectrum computed with different $\ell_{max}$ values. The parameters of the power spectrum models are taken from Table.7.
@@ 5.1 Fitting method
In this study, we concatenate the density, momentum and cross power spectrum to effectively constrain the growth rate $f\sigma_{8}$. By concatenating the models of these three spectrum, i.e. Eq.29, Eq.30 and Eq.31, we derive a comprehensive framework comprising 13 parameters in total. These encompass the growth rate $f\sigma_{8}$, the biasing parameters $b_{1}\sigma_{8},b_{2}\sigma_{8},b_{s}\sigma_{8},b_{3nl}\sigma_{8}]$ for both the density and momentum fields, as well as the velocity dispersion parameters $\sigma_{vS},\sigma_{vT}]$ for the density and momentum fields, respectively. However, in practical applications, it is not necessary to treat all 13 parameters as independent variables during model fitting with measurements, because these models don’t sensitive equally to every parameter. Through the following testing using mocks, we discover that a streamlined set of only five key parameters $\penalty 10000\ f\sigma_{8},\penalty 10000\ b_{1}\sigma_{8},\penalty 10000\ b_{2}\sigma_{8},\penalty 10000\ \sigma_{vT},\penalty 10000\ \sigma_{vS}\penalty 10000\ ]$ suffices to maintain robustness and precision in the fitting process (see Appendix E for more discussion). This leads us to adopt a unified linear biasing parameter $b_{1}\sigma_{8}$ for both the density and momentum fields, a shared second-order biasing parameter $b_{2}\sigma_{8}$ for both fields. Furthermore, we employ shared velocity dispersion parameters $\sigma_{vT},\penalty 10000\ \sigma_{vS}]$ across both the density and momentum fields.
@@ 5.2 Fitting $f\sigma_{8}$ of the mocks
The MCMC resulting parameter estimates are summarized in Table 7 and illustrated in the right panel of Fig.6. Notably, the estimated value of $f\sigma_{8}=0.468^{+0.061}_{-0.043}$, which closely aligns with the mock fiducial value of $f\sigma_{8,\rm fid}=0.466$ at effective redshift $z_{\rm eff}=0.2$. This agreement demonstrates that our fitting methodology effectively recovers the true growth rate embedded in the mocks, and that the power spectrum models perform robustly up to the non-linear scale of $k_{\text{max}}$=0.3 $h$ Mpc-1 under current conditions. In this analysis, we exclude the momentum power spectrum quadrupole, $P^{p}_{2}$, as it is heavily dominated by noise in survey data. To further assess the adaptability of our model, we also conducted fits incorporating $P^{p}_{2}$ using the mock average, where $P^{p}_{2}$ exhibits sufficient smoothness to yield a reliable fit. For a more detailed exploration, please refer to Appendix F.
The $\chi^{2}$ values reported in Table 7 are computed using the mean of the 675 catalogs, with uncertainties scaled according to a single realization. The reduced $\chi^{2}$ is 0.08 which is much sammler than one since we are fitting the mock mean with a single-mock covariance. However, when multiplying the $\chi^{2}$ by a factor of 675, the resulting value becomes excessively large. This discrepancy arises because, although the models demonstrate sufficient accuracy for our current cosmological volume, its precision is inadequate at a scale 675 times larger than our survey volume. Attempting to fit real observational data in such a vast cosmological volume would therefore yield an unsatisfactory fit. Fortunately, we are not yet at a stage where such discrepancies constitute a critical systematic concern.
@@
Table 2: The MCMC-fit cosmological parameter estimates derived from mock average. [^7] These parameter estimates are derived from the BGS and DESI-PV mocks are presented here. These estimates are obtained by fitting the average of 675 BGS mocks and 675 DESI-PV mocks. The degrees of freedom (d.o.f.) are calculated based on 120 data points and 5 free parameters.
@@
Fig.8 presents the growth rate constraints derived from fitting the power spectrum using different wave number cut-offs $k_{max}$, demonstrating that our models yield stable fits and reliable for $k_{max}<0.35\penalty 10000\ h$ Mpc-1.
Figure 8: The blue filled-squares show the estimated $f\sigma_{8}$ as a function of cut-off wave number $k_{max}$. The pink dashed- line is the fiducial value 0.466.
@@ 6.1 The systematic error of the FP data
To mitigate this bias, we have to adopt a conservative approach by restricting the momentum power spectrum analysis to $k^{p}_{\text{max}}=0.1\ h\ \text{Mpc}^{-1}$ when fitting $f\sigma_{8}$ from survey data, while retaining $k_{\text{max}}=0.3\ h\ \text{Mpc}^{-1}$ for the density and cross power spectrum. The information encoded in the momentum power spectrum is predominantly concentrated on the larger cosmological scales, specifically at wavenumbers below $k\leq 0.1\ h\ \text{Mpc}^{-1}$ (Howlett, 2019; Bautista et al., 2025). Consequently, adopting a conservative cutoff at $k^{p}_{\text{max}}=0.1\ h\ \text{Mpc}^{-1}$ enables a robust analysis of the data without sacrificing significant signal content and does not impact our main conclusion.
@@ 6.2 Result of $f\sigma_{8}$
The refined estimate of the growth rate, derived by synthesizing the fitting outcomes from correlation function (Turner et al., 2025), maximum likelihood estimation (MLE, Lai et al. 2025) and power spectrum of this paper, yields a consensus value of $f\sigma_{8}(z_{\text{eff}}=0.07)=0.450^{+0.055}_{-0.055}$, in concordance with theoretical predictions too. For a deeper exploration of data synthesis, see Bautista et al. (2025).
@@ 6.3 Result of $\gamma$
Table 3: The MCMC-fit cosmological parameter estimates derived from data. [^9] These parameter estimates are derived from the BGS and DESI-PV data are presented here. The degrees of freedom (d.o.f.) are calculated based on 98 data points and 5 free parameters.
@@ 8 Conclusions
  8. 8.
    We have explored the Galilean transformation of the power spectrum and conclude that its impact is negligible within the context of our analysis.
  9. 9.
    Finally, while most individual measurements align well with the GR prediction, a joint analysis reveals a higher $\gamma=0.63\pm 0.02$, suggesting weaker gravitational effects.
There are several promising avenues for future research building upon this work: (1) Applying our methodology to the Data Release 2 of DESI peculiar velocity survey, which is significantly more homogeneous and extensive than the Data Release 1, thereby enabling a more precise constraint on $f\sigma_{8}$; (2) While the current analysis assumes the local plane-parallel approximation, future efforts will focus on incorporating wide-angle effects into both our models and estimators; (3) As highlighted in Appendix G of this paper, the momentum power spectrum and cross power spectrum lack Galilean invariance, making them susceptible to potential systematic biases in the data. To address this limitation, we aim to transition toward measuring and modeling the pair-wise momentum power spectrum, as proposed by Chen et al. (2025); (4) We have not yet accounted for the gravitational redshift of galaxies. In forthcoming studies, we intend to extend our framework to measure and model this relativistic effect, following the approach outlined by Beutler & Di Dio (2020).
Data availability
The data used to generate the figures in this paper are available at https://zenodo.org/uploads/17672675. The public code of power spectrum measurements, power spectrum theoretical models and window function convolution are available at https://github.com/FeiQin-cosmologist/Galaxy_Power_Spectrum, and see Appendix I for more links. The examples for the utilization of the code can be found in this Jupyter Notebook https://github.com/FeiQin-cosmologist/Galaxy_Power_Spectrum/blob/main/CosmPSPy/Code/Examp_PS.ipynb
@@ Appendix A Peculiar velocity estimator
where $\sigma_{\eta}$ denotes the measurement error in the log-distance ratio $\eta$. $300^{2}$ km2 s-2 represents the intrinsic scatter in peculiar velocities caused by non-linear galactic motions. As discussed in Appendix A of Paper II, variations in the assumed value of this intrinsic scatter have minimal impact on our results. The $\langle v^{2}({\bf r})\rangle$ term for the DESI-PV random catalog is computed by Bautista et al. (2025).
@@ Appendix B Verifying $P^{\delta p}_{\ell}=-P^{\delta p*}_{\ell}$ for cross-power spectrum
The mathematical framework for converting power spectrum to correlation functions primarily draws upon Turner et al. (2025). In Fig.16, we present the model correlation functions converted from the power spectrum models presented in Section 4 with parameters from Table 7.
@@ Appendix D The multipoles of the field function
Appendix E More tests using mocks
In Section 5.2, we demonstrate the adaptability of the power spectrum models by performing fits on $P^{\delta}_{0},P^{\delta}_{2},P^{p}_{0},P^{\delta p}_{1}]$ using five free parameters $\penalty 10000\ f\sigma_{8},\penalty 10000\ b_{1}\sigma_{8},\penalty 10000\ b_{2}\sigma_{8},\penalty 10000\ \sigma_{vT},\penalty 10000\ \sigma_{vS}\penalty 10000\ ]$, which suffices to maintain robustness and precision in the fitting process. Alternatively, we can separate the biasing parameters between the density and momentum fields—i.e., adopt seven free parameters $\penalty 10000\ f\sigma_{8},\penalty 10000\ b^{\delta}_{1}\sigma_{8},\penalty 10000\ b^{\delta}_{2}\sigma_{8},\penalty 10000\ b^{p}_{1}\sigma_{8},\penalty 10000\ b^{p}_{2}\sigma_{8},\penalty 10000\ \sigma_{vT},\penalty 10000\ \sigma_{vS}\penalty 10000\ ]$—without significantly altering the resulting $f\sigma_{8}$, see Fig.17. Consequently, to maintain model simplicity and reduce parameters, we adopt shared biasing parameters for both fields.
Figure 17: The pink histograms and 2D contour plots show the fitting results obtained with $\penalty 10000\ f\sigma_{8},\penalty 10000\ b^{\delta}_{1}\sigma_{8},\penalty 10000\ b^{\delta}_{2}\sigma_{8},\penalty 10000\ b^{p}_{1}\sigma_{8},\penalty 10000\ b^{p}_{2}\sigma_{8},\penalty 10000\ \sigma_{vT},\penalty 10000\ \sigma_{vS}\penalty 10000\ ]$, while the blue counterparts are identical to those in Fig.6.
Appendix F Fitting $f\sigma_{8}$ of mocks from $P^{\delta}_{0,2}+P^{p}_{0,2}+P^{\delta p}_{1}$
@@
Table 5: The MCMC-fit cosmological parameter estimates.  [^11] Corresponding to Fig.19, the MCMC-fit cosmological parameter estimates are presented here.
@@
Appendix G The Galilean transformation of power spectrum
@@
Appendix H Deriving the estimator of power spectrum and the window funtion convolution matrix
@@
H.1 Power spectrum estimator
@@ H.3 The expression of multipoles
Appendix I Data and Code Availability
We developed a set of code for the power spectrum measurements, modeling and window function convolution. This package is called ‘CosmPS’, the GitHub link is https://github.com/FeiQin-cosmologist/Galaxy_Power_Spectrum . We have PYTHON version ‘CosmPSPy’ and FORTRAN version ‘CosmPSFt’, respectively. We also have a NERSC directory for the ‘CosmPSPy’ package, which is /global/cfs/cdirs/desi/science/td/pv/PS/CosmPSPy .
1: The ‘CosmPSPy’ package can be found in https://github.com/FeiQin-cosmologist/Galaxy_Power_Spectrum/tree/main/CosmPSPy The Jupyter Notebook for the examples of the utilization of the ‘CosmPSPy’ package (including power spectrum measurements, modeling and window function convolution) can be found in https://github.com/FeiQin-cosmologist/Galaxy_Power_Spectrum/blob/main/CosmPSPy/Code/Examp_PS.ipynb
2: The ‘CosmPSFt’ package, i.e. the FORTRAN code used to compute the measured power spectrum can be found in https://github.com/FeiQin-cosmologist/Galaxy_Power_Spectrum/tree/main/CosmPSFt
3: The hand-written draft for the derivation of the power spectrum estimator and window function convolution in Appendix H can be found in https://github.com/FeiQin-cosmologist/Galaxy_Power_Spectrum/blob/main/Math/1%20PS%20estimator.pdf
4: The Mathematica code for the derivation of the perturbation models of power spectrum in Section 4 can be found in https://github.com/FeiQin-cosmologist/Galaxy_Power_Spectrum/blob/main/Math/2%20modPS.nb or see the ‘.pdf’ version https://github.com/FeiQin-cosmologist/Galaxy_Power_Spectrum/blob/main/Math/3%20modPS.pdf
5: The PYTHON module used to compute the perturbation models of power spectrum and window function convolution can be found in https://github.com/FeiQin-cosmologist/Galaxy_Power_Spectrum/blob/main/CosmPSPy/Code/PSmodFun.py
6: The PYTHON module used to compute the measured power spectrum can be found in https://github.com/FeiQin-cosmologist/Galaxy_Power_Spectrum/blob/main/CosmPSPy/Code/PSestFun.py
7: The example PYTHON code for fiting the growth-rate $f\sigma_{8}$ from galaxy mocks can be found in https://github.com/FeiQin-cosmologist/Galaxy_Power_Spectrum/tree/main/FitExamp